OpenStudy (anonymous):

how do you find the limit of 1-cos(theta) all devide by theta, as theta move closer to zero??

OpenStudy (anonymous):

Can you use L'Hospital's rule?

OpenStudy (anonymous):

By L'Hospital's rule \[ \lim_{\theta \to 0}\frac { 1 -\cos(\theta)}{\theta}=\lim_{\theta \to 0}\frac { \sin(\theta)}{1} \] What is the answer?

OpenStudy (anonymous):

dude its seems kinda difficult please can you help me out ?

OpenStudy (anonymous):

thanks bro you are the best